OpenStudy (anonymous):

what is the solution to this equation? -7y-3=11

OpenStudy (skullpatrol):

You have posted this same question 3 times?

OpenStudy (whpalmer4):

Add and subtract to get all the y terms on one side, and all the others on the other side. What does that give you? You should then be able to divide both sides by the coefficient of y (if any) to find the value of y. Try it!

OpenStudy (anonymous):

i dont get this stuff at all

OpenStudy (skullpatrol):

What don't you understand?

OpenStudy (whpalmer4):

Well, do you understand why you can multiply or add or divide or subtract as long as you do the same thing on both sides of the equal sign? Think of it this way: you and a friend each have $10. If I give each of you $5, now you each have $10 + $5 = $15, and you still both have the same amount of money — the equation is still correct. Similarly, if I asked you to both divide your money in 3 piles and give 2 of those piles to your sister, you would each end up with $5 and so would still have the same amount of money. That's all we are doing here — doing the same thing to both sides of the equation, to make it a more convenient form.

OpenStudy (skullpatrol):

Do you know what an "inverse" operation is?

OpenStudy (skullpatrol):

Are you still there?

OpenStudy (whpalmer4):

So, -7y - 3 = 11, we want to find an expression that just has y on one side, and a number on the other. Well, let's start by getting rid of that "-3" on the left side. We do that by either subtracting it from both sides, or adding its complement (3). \[-7y - 3 + 3 = 11 + 3\]\[-7y = 14\] Now if we divide both sides by -7, we'll be left with just y on the left side, and our value of y on the right side. \[\frac{-7y}{-7}=\frac{14}{-7}\]\[y=-2\] Clear as mud?